OpenStudy (anonymous):

16x^(3)-4x^(2) -8x-15)/(4x-5)

OpenStudy (anonymous):

And you are trying to simplify this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no divide it

OpenStudy (anonymous):

the slash is a divide sign sorry

OpenStudy (anonymous):

divide or not?

OpenStudy (anonymous):

divide it not simplify

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

4x^2(4x)=16x^3 16x^3-16x^3=0 move4x^2 down 4X^2(-5)=-20x^2 -4x^2-(-20x^2)=16x^2 move -8x down 4x(4x)=16x^2, 4x(-5)=-20x -16x@-(-16x^2)=0 -8x-(-20x)=12x 3(4x)=12x , 3(-5)=-15 12x-12x=0 , -15-(-15)=0 the answer is: 4x^2+4x+3 it hard for do on key board tye,I don't know you understand or not

OpenStudy (anonymous):

yess i understood how you did it. I am working on a problem can you check the answer for me?

OpenStudy (anonymous):

I did give answer is: 4x^2+4x+3

OpenStudy (anonymous):

no i mean im doing a different problem can you check it for me once i get an answer?

OpenStudy (anonymous):

I have to go college class to study now,when I'm back I will check,bye